OpenStudy (anonymous):

How many solutions does this system of equation have? y=5x+7 and 3y-15x=18

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=5x+7\quad \Longrightarrow \quad y-5x=7 }\) \(\color{#000000 }{ \displaystyle 3y-15x=18\quad \Longrightarrow \quad y-5x=6 }\) divide the eq by 3

OpenStudy (triciaal):

divide the 2nd eqn by 3 then rearrange to the form y = now compare eqn 1 with eqn 2

OpenStudy (anonymous):

OK, thanks but only 3y -15x =18?

OpenStudy (solomonzelman):

Is it possible that \(\color{#000000 }{ \displaystyle y-5x }\) is 6 and 7 at the same time?

OpenStudy (anonymous):

BUt it seems confusing...

OpenStudy (solomonzelman):

What seems confusing?

OpenStudy (anonymous):

Like, the y-5x= 6 & 7

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y-5x }\) is jut a number (EXCEPT, that it is an unknown number - you don't know what it is equivalent to).

OpenStudy (anonymous):

So would the answer be 2 solutions? Im so confused.

OpenStudy (solomonzelman):

So, if it comes out that \(\color{#000000 }{ \displaystyle (y-5x)=7 }\) (from 1st equation) and \(\color{#000000 }{ \displaystyle (y-5x)=6}\) (from 2nd equation)

OpenStudy (solomonzelman):

would that be possible?

OpenStudy (solomonzelman):

Could same number be equal to 6 and equal to 7 (all at once)?

OpenStudy (anonymous):

No.

OpenStudy (dannyo19):

No solution

OpenStudy (anonymous):

So would that mean it needs to be 2? Or could there be more?

OpenStudy (anonymous):

Wait, no, so not 2. Yea no solution, its not possible.

OpenStudy (solomonzelman):

yes

OpenStudy (dannyo19):

You're right/

OpenStudy (anonymous):

The same number could not equal two different numbers with both,

OpenStudy (anonymous):

equations

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle (y-5x)=7 }\) (from 1st equation) \(\color{#000000 }{ \displaystyle (y-5x)=6 }\) (from 2nd equation) For this to be possible we would need to have \(\color{#000000 }{ \displaystyle 6=7 }\)

OpenStudy (anonymous):

Yes, so no solution. Thanks!

OpenStudy (solomonzelman):

and yes, must be that system will not have a solution...

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

No solution because the lines are parallel to each other. { y=7+5 x, y=6+5 x }